To prevent dehydration, intravenous (IV) solution is given to hospital patients. Commonly used are solutions containing sodium chloride and glucose. A 150,0mL IV solution contains 6.60 grams glucose (C6H12O6). What is the molarity of this solution? The molar masss of glucose is 180.16 g/mol.


1
Expert's answer
2021-02-01T03:56:34-0500

Q159928

To prevent dehydration, intravenous (IV) solution is given to hospital patients. Commonly used are solutions containing sodium chloride and glucose. A 150,0mL IV solution contains 6.60 grams glucose (C6H12O6). What is the molarity of this solution? The molar mass of glucose is 180.16 g/mol.


Solution:


We will first convert 6.60 grams of glucose to moles using the molar mass of glucose.


The volume of the solution is 150.0mL. We will convert this to 'L' by using the conversion

factor 1 L = 1000 mL.


Once we got the moles of glucose and volume of glucose solution in 'L', we can use the molarity formula and find the molarity of the solution.


Step 1: Convert 6.60 grams of glucose to moles.


Molar mass of glucose = 180.16 g/mol.


moles of glucose "= 6.60 \space g \space of \space glucose * \frac{1\space mol\space glucose }{180.16\space g\space glucose }"


= 0.036634 mol of glucose.



Step 2 : Convert 150.0 mL to 'L' by using the conversion factor 1 L = 1000 mL.


volume in 'L' = "150.0 \space mL * \frac{1 \space L}{1000 \space mL } \space = \space 0.1500 \space L"



Step 3 : Plug moles = 0.036634 mol glucose and volume = 0.1500L in the molarity formula and

find the molarity of the glucose solution.



"Molarity = \frac{moles \space of \space glucose }{volume \space of \space solution in \space 'L' }"



"Molarity = \frac{0.036634 \space mol }{0.1500\space L } = 0.2442\space mol/L"



In the question, we are given 6.60 grams of glucose in 3 significant figure, so our final answer must

also, be in 3 significant figures.


Hence the molarity of glucose is 0.242 M or 0.242 mol/L.